# DBudgetKV: Dynamic Budget in KV Cache Compression for Ensuring Optimal Performance

### v2

@@ 3.2 DBudgetKV
As the input sequence length $n$ increases, the time and space overhead for norm calculation also grows significantly by $O(n^{2})$. Hence, we next aim to reduce the computational scale from the naive norm calculation. We observe that the Frobenius norm of the last row of $A$ also correlates with the model’s performance; equivalently speaking, the attention distribution of the latest token is good enough to serve as an estimator. Formally, we utilize the last $k$ rows (latest $k$ tokens) of $A$ and reduce to a single attention vector $A^{\prime}\in\mathbb{R}^{1\times n}$. The score $s_{j}$ for a position $j\in[1,n]$ in $A^{\prime}$ is:
$s_{j}=\frac{\sum_{i=k}^{n}A_{i,j}}{\sum_{i=k}^{n}\mathbf{1}_{\{A_{i,j}\neq 0\}}}$ (1)
We then use the Frobenius norm of $A^{\prime}$, equivalently its L2 norm, as the evaluation metric, which reduces the overall complexity to $O(n)$.
Finally, for each position in the ranking order, DBudgetKV compares the difference between the Frobenius norm of the original attention matrix $A$, and the Frobenius norm of the reduced attention vector $A^{\prime}$ of which the pruned positions are set to 00. We set a universal threshold as 1%, such that when the new norm exceeds 1% difference from the original norm, it hits the stopping criterion. The KV vectors of the remaining positions are then all kept, while the preceding positions have their KV cache discarded. In this way, our method achieves dynamic KV cache pruning while targeting to maintain the full-cache performance, assisted by the norm metric.
3.3 Implemetation Details
Though our proposed pruning process is conceptually sequential, it is implemented efficiently by PyTorch’s operators, such that the stopping positions of all layers are identified directly in parallel without any looping operations. Empirical analysis in Sec. 4.4 shows that our proposed method achieves higher efficiency compared to three popular baselines.
Specifically, we reorder $A^{\prime}$ based on Token Importance Ranking $\widehat{X}$ to obtain $A^{\prime}_{sort}$, where $A^{\prime}_{sort}[0]$ corresponds to the least important token and $A^{\prime}_{sort}[-1]$ corresponds to the most important token. Next, we compute the cumulative square-sum of each element in $A^{\prime}_{sort}$:
$A^{\prime}_{cumsum}[i]=\sqrt{\sum_{k=i}^{n}|A^{\prime}_{sort}[k]|^{2}}$ (2)
@@ 3.3 Implemetation Details
such that $A^{\prime}_{cumsum}[i]$ represents the Frobenius Norm of the attention matrix after removing all tokens to the left of $i^{th}$ token. We then divide each element in $A^{\prime}_{cumsum}$ by $||A^{\prime}||_{F}$ and subtract this value from 1 to determine the difference between each new matrix and $A^{\prime}$. The torch where operation allows us to directly identify the position of the token that satisfies the condition of not exceeding a 1% difference, ultimately yielding the set of tokens for which the KV cache is called to retain. The final kv cache compression algorithm of DBudgetKV is presented in Algorithm 1. Since PyTorch allows specifying the dimension for operations, DBudgetKV can run in parallel across multiple attention heads.
@@
Additionally, we have discovered that applying DBudgetKV to the model’s 0th and 1st layers results in poor performance when the model operates under a high budget. We hypothesize that this may be due to the relatively uniform distribution of attention scores in these initial layers, leading to the premature discarding of tokens that are crucial for subsequent layers, ultimately causing the entire model’s inference to collapse. A more detailed analysis of this issue is provided in Appendix C. Based on this finding, we initiate the KV cache compression operations starting from the model’s 2nd layer. Regarding ablation studies are further conducted in Sec.4.5.
@@ 4 Experiments
4.1 Experimental Settings
Table 2: Performance of DBudgetKV and its comparison with three KV cache pruning models on 13 datasets. Bold numbers indicate the best results aside from full-cache. Italics represent the real budget utilized by DBudgetKV. The correspondence between abbreviations and their full names of tasks and datasets can be found in Appendix. Avg. calculates the mean ratio of the model’s performance using different KV cache compression methods to its performance with the full cache. All average results (except for budget) are adjusted by subtracting 1 to provide a more intuitive understanding of the effectiveness of different pruning methods.
@@ 4.1 Experimental Settings
Our experiments are conducted with three LLMs: Llama3-Instruct with size of 8B/70B, Mistral-7B-Instruct-V0.3, and Qwen2.5-Instruct with size of 7B/32B/72B. We implement DBudgetKV upon the codebase of SnapKV [^1] https://github.com/FasterDecoding/SnapKV. For the reduced attention matrix $A^{\prime}$, we set $k=1$ in practice (ablation provided in Sec. 4.5), and set $m=4$ for importance ranking.
@@
For comprehensive evaluation, we evaluate DBudgetKV using datasets of both short and long context length. For tasks of relatively short inputs, the models are assessed across mathematics, science, and commonsense reasoning on multiple datasets including GSM8K [14], GPQA [26], TheoremQA [27], TruthfulQA [28], and CoQA [29]. For long-context tasks, we adopt tasks and datasets from Longbench [30]. Appendix A provides a detailed description and statistics of these datasets along with how they are utilized in experiments.
@@
Our primary objective is to evaluate the capability of DBudgetKV to compress KV cache while striking for a full-cache performance. To this end, we compare DBudgetKV with the full KV cache performance and calculate the average compression ratio per dataset. Additionally, we select three previous KV cache compression methods with fixed budget size, including: Heavy Hitter Oracle (H2O) [8], StreamingLLM (SLM) [10], and SnapKV [18]. We compare with three budgets for these three methods at 90%, 50% and 20% compression ratio respectively.
Apart from the main experiments, we further evaluate various design factors of DBudgetKV via ablation studies, such as different strategies of token ranking importance, the selection of $k$ for the attention matrix, and the necessity to retain bottom layers. Finally, we also provide efficiency analysis on the time and space overhead comparing against other KV cache pruning methods in Sec. 4.4.
@@ 4.2 Main Results
The main results are shown in Table 2 according to our evaluation protocol, where DBudgetKV demonstrates strong advantages with robust performance across diverse tasks and models:
Table 3: Performance of LLMs with scales of 70B, 32B, and 72B across five datasets using DBudgetKV.
GSM8K CoQA NQA Musique QMSum
Llama3-70B-Instruct
Full 89.69 60.36 27.15 29.31 22.52
Oursk=1 89.76 60.38 26.94 28.88 22.30
Budget 93.3% 77.2% 58.7% 60.8% 52.9%
Qwen2.5-32B-Instruct
Full 91.36 58.03 24.78 40.04 22.84
Ours 91.81 57.29 22.65 40.54 22.52
Budget 91.6% 85.0% 68.2% 74.4% 76.8%
Qwen2.5-72B-Instruct
Full 90.22 54.14 24.36 42.13 23.93
Ours 90.30 54.19 24.10 41.70 23.31
Budget 90.9% 87.6% 63.4% 65.0% 68.6%
@@
• DBudgetKV is able to fulfill our objective, capable of performing near-lossless dynamic compression across different models, varying input lengths, and diverse task types. Interestingly, with Llama3-8B and Qwen2.5-7B, DBudgetKV even surpasses the full-cache performance by 0.12% and 2.63% respectively, utilizing an average of 63.68% and 76.02% KV cache. With Mistral, DBudgetKV also manages to achieve nearly 20% KV cache compression with only a 1.5% performance reduction. These results indicate that our proposed method can be suited for real-world scenarios with no bother by domain-specific budget thresholds. In stark contrast, previous methods achieve a consistent full-cache performance only when manually determined a high budget ratio, e.g. 90%. However, when the budget is reduced, e.g. 50% and 20%, the degradation can become severe on certain datasets, distinct from DBudgetKV that automatically adjusts the pruning to always prioritize full performance.
@@
• DBudgetKV naturally reflects the difficulty of the generation task. As in Table 2, the dynamic budget ratio is high on Math&Science datasets (over 90%), while much lower on QA or Summarization datasets (as low as 15%). This observation is in line with our intuition, where inference on concise but hard tasks, such as math problems, requires more context and more precise calculation, resulting in higher budget allocation. Thus, our method design serves beyond for inference efficiency, but could also be leveraged for input analysis in a broader scope.
• Besides the dynamic compression, DBudgetKV also outperforms the three 90%-budget baselines, while itself uses less than 90% budget, and achieves the highest score in 20 out of 39 comparisons. This suggests that by taking into account both the position bias and attention scores, our method accompolishes more effective pruning than the baselines that only consider one aspect, highlighting the importance of integrating both dimensions for effective KV cache compression.
It is worth noting that the scope of this work is not to propose yet another KV cache pruning method. Instead, we identify the common limitation of previous KV pruning methods, and call for optimal KV pruning without a pre-defined budget. Consequently, the consistent performance across domains after KV pruning is the ultimate goal of our method, instead of achieving the highest compression ratio on certain datasets. Under our objective, we leave the further improvement on compression ratio as a promising future research direction.
4.3 DBudgetKV’s Generalization Ability
We next conduct experiments to examine whether our design and hyper-parameters can be generalized to LLMs of larger sizes. On five datasets in Table 3, DBudgetKV with Llama3-70B and Qwen2.5-32B/72B demonstrates consistent near full-cache performance with the same DBudgetKV setting. Notably, the averaged compression ratio increases on datasets of longer context, achieving nearly a 50% compression. But most importantly, they are able to maintain lossless performance across inputs, validating the robustness and versatility for its general applicability.
4.4 Efficiency Analysis
@@
Table 5: Performance comparison for ablation studies in Section 4.5, with different $k$ values for attention matrix reduction, ranking methods and universal thresholds.
@@
Posk=1 Performance with Different Threshold $t$
${t=0.1\%}$ 76.35 52.84 22.73 17.04 21.90
Budget 99.2% 97.0% 77.2% 85.0% 73.9%
${t=10\%}$ 26.23 49.69 18.35 15.16 18.81
Budget 55.8% 37.2% 15.27% 18.5% 6.5%
In this section, a quantitative study is conducted on whether DBudgetKV improves inference time, apart from the space reduction from KV cache pruning. We compare the time usage on six datasets with Llama3-8B/70B using DBudgetKV, along with three baselines with the 50% budget setting. Table 4 reports the average generation time after the prefilling stage of each sample (Overall), as well as the time needed to complete pruning (Prune) before the generation.
From the results, DBudgetKV achieves the best performance in 8 out of 12 comparisons of overall generation time, and achieves the best average time for both 8B and 70B LLMs, suggesting that although DBudgetKV requires a slightly longer time for pruning itself compared to three other approaches, the total generation speed of DBudgetKV is evidently advantageous. It also performs consistently across models of different scales, underscoring its time efficiency.
4.5 Ablation Study
The ablation study is conducted to explore the impact of various configurations of DBudgetKV. We select Llama-3-8B-Instruct and perform experiments across five datasets presented in Table 5. Appendix B presents additional results and analysis from the ablation study.
@@ 4.5 Ablation Study
The reduced attention matrix $A^{\prime}$ in Sec. 3.2 aggregates attention scores from the last $k$ positions. The upper part of Table 5 illustrates the model’s performance and the actual budget when setting $k=1$, $1\%n$, $5\%n$, and $10\%n$ ($n$ being the number of input tokens). It is clear that setting $k$ as $1$ achieves a significantly reduced budget, thus a higher compression ratio, with almost no change in performance compared to $1\%n$ and $5\%n$. On the other hand, while $10\%n$ can compress more KV cache, it fails to maintain performance (for instance, on NarrativeQA, the former achieves a performance of 9.74 using 32% of the budget, whereas the latter scores 21.44 using 48.7% of the budget). What’s even better is that since $k=1$ requires the least amount of computation, relying solely on the scores from the last token, the complexity of obtaining $A^{\prime}$ becomes $O(1)$, independent of the sequence length. The advantages of both high efficacy and low overhead make $k=1$ a solid design choice for calculating the norm metric.
Importance Ranking Strategies
We investigate DBudgetKV potentials through ranking by each token’s average attention score received from other tokens, similar to previous approaches such as H2O. The results are shown in the middle of Table 5: under the same experimental settings, the attention-based ranking struggles to maintain the model’s performance. Therefore, the position-based strategy is more suited for our goal.
Table 6: Performance comparison between DBudgetKV and StreamingLLM with the first two layers frozen (SLMF) under difference KV cache budgets ($0.9$, $0.5$ and $0.2$).
GSM8K GPQA CoQA NrtvQA QMSum TriviaQA
Full 75.28 29.02 52.74 24.06 22.27 70.26
Oursk=1 76.50 30.13 52.86 23.44 21.95 64.24
Budget 93.2% 86.7% 81.5% 48.7% 15.0% 41.0%
SLM0.9 72.78 28.79 52.83 23.75 22.48 69.97
SLMF0.9 76.42 29.24 52.89 22.95 22.47 69.10
SLM0.5 3.41 6.03 44.68 20.94 21.80 67.79
SLMF0.5 4.55 3.13 44.59 21.31 21.88 68.04
SLM0.2 3.21 3.35 40.11 20.12 21.11 66.17
SLMF0.2 1.21 1.12 38.44 19.52 20.61 65.69
Attention Norm Threshold
As we adopt the universal threshold as 1% for the attention norm difference, we comprehensively study the effects of smaller or larger thresholds, as shown in the bottom part of Table 5. Intuitively, a larger threshold allows for a higher compression ratio, while a smaller threshold does the opposite. The conclusion is clear that when the threshold is set to a smaller value, such as 0.1%, the average budget increases as expected, yet the model’s performance sees little improvement. Conversely, when the threshold is set to 10%, it prunes more KV cache, but the model’s performance significantly deteriorates. Thus, we deem 1% as a reasonable threshold for universal uses across models and tasks.
Comparison with StreamingLLM

### v3

@@ 2.2 ReFreeKV
As the input sequence length $n$ increases, the time and space overhead for norm calculation on the attention matrix grows by $O(n^{2})$. To reduce the computational scale, we seek to use an approximate norm calculation by $O(n)$. Instead of using the full attention $A$, we reduce $A$ by taking the average of its last $k$ rows to a single attention vector $A^{\prime}\in\mathbb{R}^{1\times n}$. The score $s_{i}$ for a position $i\in[1,n]$ in $A^{\prime}$ is denoted as:
$s_{i}=\frac{\sum_{j=k}^{n}A_{i,j}}{\sum_{j=k}^{n}\mathbf{1}_{\{A_{i,j}\neq 0\}}}$ (2)
Logistics
The rationale behind ReFreeKV ’s two-stage process is that finding the optimal pruning positions in a sequence is a combinatorial problem. By combining ranking with a linear search, the problem becomes tractable. The entire procedure is further optimized through parallel operations, as described as follows.
Figure 2: Performance trends of Llama3-8B, Mistral-7B, and Qwen2.5-7B across varying Uni-Metric thresholds. The x-axis represents the threshold percentage (0.1% to 5%), and the y-axis denotes the performance score normalized by the full-cache performance.
2.3 Implementation Details
As ReFreeKV is conceptually a sequential search process, the design of ReFreeKV allows efficient implementation by PyTorch’s operators, such that the stopping positions of all layers are directly identified in parallel without explicit looping operations. Latency and throughput analyses in Section 3.3 and Appendix F demonstrate that ReFreeKV matches prior popular KV pruning approaches, with negligible latency overhead and improved throughput compared to the baselines.
Specifically, the pruning position $i_{\text{prune}}$ can be determined directly by the combination of Torch cumulative-sum and where operators. We compute the cumulative square-sum of each element in $A^{\prime}_{\text{rank}}$, obtained after the initial ranking on $A^{\prime}$:
[…]
@@ 2.3 Implementation Details
such that $A^{\prime}_{cumsum}[i]$ represents the Frobenius Norm of the attention matrix after removing all cache to the right of position $i$. We then divide $\widetilde{A^{\prime}_{i}}||_{F}$ by $A^{\prime}||_{F}$ to determine the norm difference as in Eq (1). The torch where operation allows us to directly identify the leftmost position that satisfies the $1\%$ universal threshold, ultimately yielding the set of positions for which the KV cache is called to retain. The overall pruning process of ReFreeKV is further presented in Algorithm 1.
@@
In our experiments, we discovered that the LLM’s first two Transformers have a relatively uniform attention distribution, usually requiring to retain most of the cache positions. For simplicity and robustness, we always retain all KV cache of the first two layers in our implementation. More studies on bottom layers are further provided in Appendix B.
Running at Scale
Supporting batch sizes >1>1, ReFreeKV is able to perform the entire pruning process for each sample in parallel. To achieve this, we pad the shorter cache segments and update the attention masks accordingly, allowing the LLM to ignore the padded KV positions. The padding operation has a negligible impact on overall performance. Meanwhile, in popular LLM inference engines, e.g. vLLM kwon2023efficient, it is possible to allocate separate KV cache size for each sample, which aligns well with ReFreeKV. Integration with LLM inference engines such as vLLM is under planning for future development.
@@ 3 Experiments
PyramidKV0.2 9.70 8.57 9.60 26.42 51.19 26.42 39.48 39.54 25.73 21.02 21.65 85.31 1.76 -23.47%
CAKE0.2 7.58 9.11 5.56 23.66 49.19 30.20 43.04 46.35 29.00 23.33 22.94 85.45 2.89 -16.98%
Table 2: Performance of ReFreeKV and its comparison with five KV pruning methods on 13 datasets. Bold numbers indicate the best results aside from full-cache. Italics represent the real budget utilized by ReFreeKV. The correspondence between abbreviations and their full names of datasets can be found in Appendix. Avg. calculates the mean ratio of the model’s performance using different KV cache compression methods to its performance with the full cache. All average results (except for budget) are adjusted by subtracting 1 to provide a more intuitive understanding of the effectiveness of different methods.
3.1 Experimental Settings
@@
Our experiments are conducted with three LLM families of different model sizes: Llama3-Instruct with size of 8B/70B, Mistral-7B-Instruct-V0.3, and Qwen2.5-Instruct with size of 7B/32B/72B. We implement our ReFreeKV upon the released codebase of SnapKV***https://github.com/FasterDecoding/SnapKV. For the reduced attention matrix $A^{\prime}$, we set $k=1$ in practice (ablation provided in Section 3.4), and set $m=4$ for the initial cache ranking stage.
@@
For comprehensive evaluation, we evaluate ReFreeKV on datasets of both short and long context length of different domains, including mathematics, science, and commonsense reasoning on GSM8K DBLP:journals/corr/abs-2110-14168, GPQA DBLP:journals/corr/abs-2311-12022, TheoremQA DBLP:conf/emnlp/ChenYKLWMXWX23, TruthfulQA DBLP:conf/acl/LinHE22, and CoQA DBLP:journals/tacl/ReddyCM19. We also include tasks from Longbench DBLP:conf/acl/BaiLZL0HDLZHDTL24 with 8 datasets spanning document comprehension, summarization and coding. Appendix D provides a detailed description and statistics for all 13 datasets, along with how they are utilized in our experiments.
@@
Our evaluation primarily assesses ReFreeKV ’s ability to preserve full-cache performance with its automatic pruning budgets. Accordingly, we compare ReFreeKV against its full-cache scores and also report the average compression ratio for each dataset.
Additionally, we compare with five prior KV cache pruning methods with varied budget sizes, including: Heavy Hitter Oracle (H2O) DBLP:conf/nips/Zhang00CZC0TRBW23, StreamingLLM (SLM) DBLP:conf/iclr/XiaoTCHL24, SnapKV DBLP:conf/nips/LiHYVLYCLC24, PyramidKV DBLP:journals/corr/abs-2406-02069 and CAKE DBLP:conf/iclr/QinCLHFCLL25. By evaluating performance consistency under fixed KV cache budgets of 90%, 50%, and 20%, we further illustrate the limitations arising from the input-specific threshold dependence.
Lastly, we also include a concurrent work Twilight DBLP:journals/corr/abs-2502-02770, which does not rely on a budget threshold but employs a top-$p$-inspired metric for adaptive token selection. It is worth noting that the $p$ value remains a hyperparameter to be tuned across models and inputs.
@@ 3.2 Main Results
The main experimental results are shown in Table 2. Due to space limitations, the results for the 50% budget size are placed in Appendix C. For comparison with Twilight, we separately report the results in Table 4 due to its different hyperparameter type. Based on Table 2, we can draw the following observations.
@@
• ReFreeKV is able to fulfill our objective, capable of performing near-lossless dynamic compression across different models, varying input lengths, and diverse task types. Interestingly, with Llama3-8B and Qwen2.5-7B, ReFreeKV even surpasses the full-cache performance by 0.12% and 2.63% respectively, utilizing an average of 63.68% and 76.02% KV cache. With Mistral, ReFreeKV also manages to achieve near 15% compression with a relatively small 1.5% performance reduction. These results indicate that our proposed method can be suited for real-world scenarios with no bother by input-specific budget thresholds.
• In stark contrast, previous methods achieve a consistent full-cache performance only when manually determined a high budget ratio, e.g. 90%. However, when the budget is reduced, e.g. 50% and 20%, the degradation can become severe on certain datasets, distinct from ReFreeKV that automatically adjusts the pruning to always prioritize full-cache performance. Theoretically, one could tune the budget for each dataset that achieves minimal degradation, but this is generally infeasible for real-world open-domain instructions.
@@
• ReFreeKV naturally reflects the difficulty of the generation task. As in Table 2, the dynamic budget ratio is high on Math&Science datasets (over 90%), while much lower on QA or Summarization datasets (as low as 15%). This observation is in line with our intuition, where inference on concise but hard tasks, such as math problems, requires more context and more precise calculation, resulting in higher budget allocation. From this aspect, our method design serves beyond for memory efficiency, but could be potentially leveraged for input analysis in a broader scope.
• Besides the dynamic compression, ReFreeKV also outperforms the three 90%-budget baselines, while itself uses less than 90% budget. Though, It is worth noting that the goal of this work is not to propose yet another KV cache pruning method that targets the best possible compression ratio under specific conditions. Instead, we seek to lift the threshold constraints and advocate for robust KV pruning that generalizes to arbitrary inputs. Through our proposed objective, we hope to further advance research in this direction.
@@
Methods GSM8K NQA Qasper 2WQA Musique
Llama3-8B 75.28 24.06 43.91 35.33 14.77
+ Oursk=1 76.50 23.44 37.38 36.21 15.96
+ Twip=0.95 76.04 23.37 43.08 36.18 14.92
Mistral-7B 33.36 28.74 37.80 33.87 22.88
+ Oursk=1 31.54 27.20 38.93 33.46 22.15
+ Twip=0.85 30.33 27.17 38.90 33.23 22.92
Table 4: Performance comparison between ReFreeKV and Twilight across five datasets using Llama3-8B-Instruct and Mistral-7B-V0.3. Bold numbers indicate the best results aside from full-cache.
As separately reported in Table 4, we compare ReFreeKV with Twilight’s reported performance across the GSM8K, NarrativeQA, 2WikiMQA, and Musique datasets. Regarding the hyperparameter for Twilight, we adopt their reported optimal $p$ value: $p=0.95$ for Llama3-8B and $p=0.85$ for Mistral. As shown in the results, ReFreeKV achieves performance comparable to Twilight on both models. Notably, both methods frequently surpass the full-cache baseline across these datasets. Though, as Twilight requires different optimal $p$ values for each model, tuning may still be necessary. Nonetheless, both ReFreeKV and Twilight share the same principle of adaptive pruning, which we believe can further advance research in this area.
3.3 Efficiency Analysis
In this section, we conduct a quantitative study on the latency of ReFreeKV and the overall impact on the inference time, apart from the space reduction from KV cache pruning. We compare the time usage on six datasets with Llama3-8B/70B using ReFreeKV, along with three baselines with the 50% budget setting. Table 3 reports the time latency needed to apply pruning (Prune), as well as the average generation time after the prefilling stage of each sample (Overall).
From the results, the latency by ReFreeKV is on par with prior pruning methods. Notably, ReFreeKV achieves the best performance in 8 out of 12 comparisons of overall generation time, suggesting that the total generation speed of ReFreeKV is advantageous. The trend also remains consistent across different model sizes, underscoring the efficiency aspect of ReFreeKV.
For batch process, we further conduct analysis on throughput with different batch sizes. The full details of which are provided in Appendix F. The results confirm that ReFreeKV improves throughput upon the naive generation by 10-20%, especially it retains performance edges and robustness with increasing batch sizes.
3.4 Ablation Studies
We conduct ablation studies to investigate the impact of various configurations of ReFreeKV. We use Llama-3-8B-Instruct and perform experiments across five datasets, with results presented in Table 5. Appendix E provides additional results and analysis from our ablation studies.
Figure 3: Performance vs. Efficiency Trade-off. (a) Performance retention across five datasets (solid lines). (b) Computational budget consumption (dashed lines) relative to the dense baseline. The shared legend indicates the datasets. Results show that setting the universal threshold to $1\%$ could well balance between performance and memory budget, as it maintains robust full-cache performance while substantially reducing KV cache.
@@ 3.4 Ablation Studies
Table 5: Performance comparison for ablation studies in Section 3.4, with different $k$ values for attention matrix reduction and initial ranking alternatives.
@@
The reduced attention matrix $A^{\prime}$ in Section 2.2 aggregates attention scores from the last $k$ rows. The upper part of Table 5 illustrates the model’s performance and the actual budget when setting $k=1$, $1\%n$, $5\%n$, and $10\%n$ ($n$ being the number of input tokens). It is clear that setting $k$ as $1$ achieves a significantly reduced budget, thus a higher compression ratio, with almost no change in performance compared to $1\%n$ and $5\%n$. On the other hand, while $10\%n$ can compress more KV cache, it fails to maintain performance (for instance, on NarrativeQA, the former achieves a performance of 9.74 using 32% of the budget, whereas the latter scores 21.44 using 48.7% of the budget). What’s even better is that since $k=1$ requires the least amount of computation, relying solely on the scores from the last row, the complexity of obtaining $A^{\prime}$ becomes $O(1)$, independent of the sequence length. The advantages of both high efficacy and low overhead make $k=1$ a solid design choice for calculating the norm metric.
Initial Ranking Strategies
Apart from the position-based ranking described in Section 2.2, we also investigate other alternatives, such as ranking by each token’s average attention score received from other tokens, similar to previous approaches such as H2O. The results are shown in the bottom of Table 5: under the same experimental settings, the attention-based ranking struggles to maintain a robust final performance. Empirical validation supports that position ranking is an appealing choice, offering both superior efficacy and efficiency.
The Universal Threshold
As we adopt the universal threshold as 1% for the attention norm difference, we comprehensively study the effects of smaller or larger thresholds, as shown in Figure 3. Intuitively, a larger threshold allows for a higher compression ratio but potentially more degradation, while a smaller threshold does the opposite. It is clear that when the threshold is set to a smaller value, such as 0.1%, the average budget increases as expected, yet the model’s performance sees little improvement. Conversely, when the threshold is set to 10%, it prunes more KV cache, but the model’s performance significantly deteriorates. Thus, we deem 1% as a reasonable and robust threshold for universal uses across models and tasks.
Generalization
We further conduct experiments to examine whether our design and hyperparameters can be generalized to LLMs of more sizes. On five datasets in Appendix Table 12, ReFreeKV with Llama3-70B and Qwen2.5-32B/72B demonstrates consistent near full-cache performance with the same ReFreeKV setting. Notably, the averaged compression ratio increases on datasets of longer context, achieving nearly a 50% compression.